OpenStudy (jojo12g):

will fan and medal please help A rectangular prism and a rectangular pyramid have the same length, width, and height. What is true about the volumes of the rectangular prism and the rectangular pyramid? A Not enough information is given to determine a relationship between the volumes. B The rectangular prism has a volume one-third the volume of the rectangular pyramid. C The rectangular prism has a volume three times the volume of the rectangular pyramid. D The rectangular pyramid has a volume three times the volume of the rectangular prism.

OpenStudy (jojo12g):

@Will.H

OpenStudy (xapproachesinfinity):

you want to think about the volumes of each and compare them

OpenStudy (xapproachesinfinity):

rectangular prism is just WLH Pyramid though will differ

OpenStudy (jojo12g):

i think D

OpenStudy (xapproachesinfinity):

well IDK we need to justify that

OpenStudy (xapproachesinfinity):

rectangular pyramid has volume V=1/3 WLH

OpenStudy (xapproachesinfinity):

They told us they both have the same L, W, H

OpenStudy (xapproachesinfinity):

D is wrong as you can see it say rectangular pyramid is 3 time the volume of the prism which not true since we have V=1/3 LWH

OpenStudy (will.h):

Rectangular prism volume formula = L*W*H Rectangular pyramid volume = (1/2)A*C*H Now plug a congruent values for example l=2 and w= 1 and H= 2

OpenStudy (xapproachesinfinity):

so it actually a third of the prism

OpenStudy (xapproachesinfinity):

it is 1/3 not 1/2 dear

OpenStudy (will.h):

You sure?

OpenStudy (xapproachesinfinity):

yeah 100% sure :)

OpenStudy (will.h):

Thanks for the correction :)

OpenStudy (xapproachesinfinity):

no problem we are here to correct each other and learn

OpenStudy (will.h):

Fanned you :)

OpenStudy (xapproachesinfinity):

answer should be at your disposal now if D is incorrect

OpenStudy (xapproachesinfinity):

thanks :)

OpenStudy (jojo12g):

so is it B

OpenStudy (mathstudent55):